Appendix A, Problem A/011 Multistep Determine the moment of inertia of the shaded area about the x-axis. у - 15 5 16.0 ParPart 2 x Incorrect Calculate the moment of inertia about the x-axis. dx 5 16.04 v ly=b-kx² 15 5757.29 4 Answer: Ix
Answer #1

solution: - at x=0 at x= 15, y zo 2. y=16 y = buka2 (16=b) Hostel -y=b-ka2 0= 16- K(15) Jet Lli -15 K=0.071 now, moment ofnow moment of inertia of shaded about ac- quis is: areg I= I, - Is I = 46305-27 807 77 Ang I=18497.25 int